OpenStudy (bloomlocke367):

What are the asymptotes of the hyperbola with the equation 9y^2-4x^2=36? Graph the hyperbola.

OpenStudy (bloomlocke367):

I already put it in standard form

OpenStudy (anonymous):

pretty much like the last one right? btw no one likes these conic section questions i noticed they are not that hard

OpenStudy (bloomlocke367):

and I have an idea what it looks like.. what are asymptotes?

OpenStudy (anonymous):

\[y=\pm\frac{a}{b}x\]

OpenStudy (anonymous):

that is of course the form if it is \[\frac{y^2}{a^2}-\frac{x^2}{b^2}=1\]

OpenStudy (bloomlocke367):

so... I'm kinda confused.

OpenStudy (anonymous):

what general form did you get?

OpenStudy (anonymous):

after dividing by 36 that is

OpenStudy (bloomlocke367):

\[\frac{ y^2 }{ 4}+\frac{ x^2 }{ 9 }=1\]

OpenStudy (anonymous):

with a minus sign not a plus sign right?

OpenStudy (anonymous):

that means \(a=2,b=3\)

OpenStudy (anonymous):

asymptotes are \[y=\frac{2}{3}x\] and \[y=-\frac{2}{3}x\]

OpenStudy (bloomlocke367):

yes, my bad

OpenStudy (bloomlocke367):

ohhh.... okay :) so it's the slope?

OpenStudy (bloomlocke367):

and the y-intercepts are \[(0,\pm2)\] right?

OpenStudy (bloomlocke367):

@satellite73